# From User Preferences to Base Score Extraction Functions in Gradual Argumentation

### v1

@@ 5. Evaluation
This section presents a theoretical validation of the concrete base score extraction functions based on how they satisfy the theoretical properties we defined. We undertake an experimental evaluation, showing the pros and cons of each base score extraction function. Finally, we give an intuition of which gradual semantics to select depending on the decision-making context.
@@ 5.1. Theoretical Analysis
First, we show under which conditions our proposed base score extraction functions satisfy the axioms and properties presented. This is summarised in Table 1. The proofs can be found in Appendix A.
@@
Setting the constants $\Delta>0$ and $\delta>0$ satisfies Axiom 1 for both base score extraction functions.
Preference ordering $\succeq_{p}$ Design Choices QE Strengths EB Strengths DF Strengths
@@
Table 2. Examples of preference orderings with different design choices for the base score extraction function 1. The value in bold between $\sigma_{slow}$ and $\sigma_{fast}$ is the system output. The (-) symbol indicates when a design choice is not needed, which is the case where all arguments are equally preferred, and their base scores are set to 0.5.
@@
Setting $\Delta>\delta>0$ satisfies Axiom 2 for both base score extraction functions.
Since the base score extraction functions are based on the distance between arguments given an ordering, if that ordering is changed, the base scores’ ordering will also change.
@@
Both base score extraction functions satisfy Axiom 3.
@@
To have regular preference relations, the variables of the algorithm to compute the distance ($\Delta,\delta$) must be constant.
@@
Both base score extraction functions satisfy Base Score Preference Stability.
@@
Given a base score extraction function $\nu$, where $\nu$ satisfies Axiom 1, and a semantics $\sigma$, where $\sigma$ satisfies monotonicity and balance, for any arguments $a$, $b$ where $a\succ b$ and $a$ and $b$ have equivalent supporters and attackers, then $\sigma(a)>\sigma(b)$.
@@ 5.2. Experimentation
To explore the robustness of our approach, we generated 30,000 random samples of preference orderings and design choices, recording the option selected by three gradual semantics models. These models are: the Quadratic Energy (QE) Potyka_kr18, the Euler-Based (EB) Amgoud_IJAR18, and the DF-QuAD (DF) Rago_kr16, which are among the most used in the literature. The decision arguments’ base scores are set to 0.5. We computed pairwise agreement and Cohen’s Kappa coefficient ($\kappa$) to assess consistency between models. Under the centralisation property, the pairwise agreement values were QEEB=0.98QE-EB=0.98, QEDF=0.85QE-DF=0.85, and EBDF=0.84EB-DF=0.84. The corresponding Cohen’s Kappa scores were QEEB=0.96QE-EB=0.96, QEDF=0.65QE-DF=0.65, and EBDF=0.62EB-DF=0.62. Without centralisation, the pairwise agreement values were QEEB=0.96QE-EB=0.96, QEDF=0.81QE-DF=0.81, and EBDF=0.81EB-DF=0.81, with Cohen’s Kappa scores of QEEB=0.92QE-EB=0.92, QEDF=0.52QE-DF=0.52, and EBDF=0.52EB-DF=0.52. These results show a strong alignment between QE and EB, while DF-QuAD diverges more, particularly when base score ranges are extreme (e.g., $\tau(e)\approx 1$, $\tau(e)\approx 0$). This difference shows the structural properties of the methods discussed in Sec. 5.3.
To further analyse the influence of user preferences and design parameters, Table 2 reports examples of preference orderings combined with different design choices, evaluated under the three gradual semantics. The preference relations are considered regular to avoid complexity in experimentation and noise in the results.
@@ 5.3. Discussion on the Gradual Semantics
Figure 2. Influence of an argument’s base score and its supporter or attacker on the final argument strength under different gradual semantics. Each subplot corresponds to a distinct semantics: DF-QuAD (left), Quadratic Energy Model (centre), and the Euler-based Semantics (right). The x-axis represents the base score of the influenced argument, and the y-axis its resulting final strength. Blue lines denote when the influencing argument is a supporter (+), while red lines denote when it is an attacker (-). Each line has assigned its aggregation ($\alpha$), with colour intensity indicating its strength.
@@
Figure 2 shows the influence of a single attacker or supporter (the influencer), and an argument’s base score (the influenced) for its final strength. The DF-QuAD semantics use product aggregation with a linear influence function, which causes the influencer’s base score to dominate the final outcome. When an attacking (or supporting) argument has a base score close to 1, it almost completely drives the influenced argument’s strength toward 0 (or 1), regardless of the influenced base score. This strong reactivity can be desirable in safety-critical scenarios, where highly preferred arguments should decisively determine the outcome, but it may reduce robustness in settings with uncertain or noisy preferences. In contrast, the QE combines sum aggregation with a 2-Max influence function. Here, both the influencer and the influenced arguments contribute comparably to the final strength, leading to smoother transitions and more gradual adaptation to preference changes (e.g., given an influencer with a base score of 1, the influenced’s final strength will be $\sigma=0.5\tau+0.5$ if it receives a support, and $\sigma=0.5\tau$ if it receives an attack). The QE, therefore, offers a good compromise between responsiveness and stability, which makes it particularly suitable for explainable decision-support systems that require consistent reasoning traces. Finally, the EB is the most conservative, using sum aggregation with an Euler-based influence function. When an argument’s base score is close to the extremes (0 or 1), its final strength remains nearly fixed at those values, regardless of new supports or attacks. This behaviour preserves high-confidence arguments, making EB appropriate when users insist on arguments being non-negotiable or when decisions must respect rigid value priorities.
@@ 6. Related Work
Gradual argumentation is highly reliant on the arguments’ base scores, and its literature has unveiled different ways of getting those, depending on the context. First, in Baroni_AC15, the authors presented the QuAD framework, intending to support debates on design alternatives. Since those scenarios do not involve a large number of voters to compute a base score, the recommended approach is to have an expert assess the base score, or obtain it from a given set of criteria and their importance. As an extension, in Rago_icppmas17, they presented QuAD-V, which allows users to vote in favour of or against arguments. The votes are used to compute a base score for each argument according to multiple users. A similar procedure is done in Tarle_aamas22. In another work Cocarascu_aamas19 in which a QBAF is generated from movie reviews to extract the most relevant information, the base scores are obtained from an aggregation of critics’ votes. Meanwhile, within structured argumentation, the work in Tamani_icfs14 is focused on ASPIC argumentation frameworks with fuzzy set theory and uses the expertise knowledge to adjudicate the importance of an argument in an environment, similarly to a base score. Another use of QBAFs is to apply their methods for explainable AI, mainly transforming (deep) neural networks into QBAFs. For instance, in Albini_arxiv20, the base scores of the arguments are obtained from the attackers and supporters and activations in the network, similarly to the work in Potyka_aaai21. Differently from the previous approaches, the work from Battaglia_icsum24 sets the base scores based on users’ preferences, similarly to this work. The preferences are integrated to personalise decision support. In a nutshell, the initial value of the base scores is arbitrarily set, and then a discount factor is applied to the less preferred arguments. This work proposes a method to use preferences to determine the initial value of the base scores.
In summary, two main limitations are found in existing work. First, most approaches do not consider the preferences of a single user to set the base scores, making it difficult to personalise argumentation-based decision-making systems. Second, when preferences are considered, the base scores are arbitrarily initialised and then adapted based on the preferences. In this work, we address those research gaps.
@@ 7. Conclusions and Future Work
This work introduces base score extraction functions, a principled method to derive argument base scores from user use preference orderings in gradual argumentation. We defined a set of axioms and desirable properties, proposed two concrete base score extraction functions with tunable design choices, and incorporated non-linear preference intensity to better capture human reasoning. Theoretical analysis confirmed that the functions satisfy coherence and regularity properties, while experiments in an assistive robotics scenario showed that preference ordering dominated the decision outcome and design choices primarily modulate sensitivity.
Future directions include the fine-tuning or definition of the different design choices given the context and user feedback, allowing partial orderings in preferences in case some arguments cannot be compared Visser_wtafa11, and computing the aggregation when using this approach when considering more than one user’s preferences in the decision-making Rago_AI25. A philosophical question which should be commented on is whether it is realistic to consider the scenario in which the most preferred argument supports the least preferred argument, which is not considered in this work. Finally, the extension of this work into value-based argumentation could provide an approach for more value-aligned decision-making Atkinson_AL21; Bodanza_AC23. This could be potentially achieved by grouping the arguments into sets of values and ordering the values according to the user’s preferences.
@@ Appendix A Proofs
Let $a,b\in\mathcal{X}$ be two adjacent arguments in the ordering, such that $a\succ b$ in the preferencing ordering. According to the Definition 9 and Algorithm 1, the distance assigned to $a$ and $b$ satisfies $d_{b}=d_{a}+\delta$. Where $\delta>0$ for a regular preference ($\succ$) and $\Delta$¿0 for a much greater preference ($\succ\!\succ$). Both proposed base score extraction functions $\nu_{1}$ and $\nu_{2}$ are strictly decreasing with respect to $d$. Therefore, if $d_{a}<d_{b}$, it follows that $\nu(D,d_{a})>\nu(D,d_{b})$. This implies that the base score of the preferred argument $a$ is greater than that of $b$, i.e., $\tau(a)>\tau(b)$. Hence, setting $\delta>0$ (and similarly $\Delta>0$) ensures that both base score extraction functions satisfy Axiom 1.
@@
$d_{b}=d_{a}+\Delta,\qquad d_{c}=d_{b}+\delta,\qquad d_{d}=d_{c},$
@@
which satisfies Axiom 2. Therefore, setting $\Delta>\delta>0$ guarantees that both $\nu_{1}$ and $\nu_{2}$ respect the intended magnitude differences between much greater and greater preferences.
@@
Let $\nu$ be a base score extraction function applied to a set of arguments $X$ with a preference ordering $\succeq_{p}$. By definition, $\nu_{p}(X)$ assigns base scores according to the relative ordering and distances among the arguments.
Suppose now that the preference ordering changes to a non-isomorphic ordering $\succeq_{p^{\prime}}$. By Definition 1, two preference orderings are isomorphic if and only if there exists a bijection $f:X\to X^{\prime}$ such that, for all $i,j\in X$,
$i\succeq_{p}j\;\Leftrightarrow\;f(i)\succeq_{p^{\prime}}f(j).$
If $\succeq_{p}$ and $\succeq_{p^{\prime}}$ are not isomorphic, then at least one pair of arguments $(i,j)$ exists such that their preference relations differ, i.e.,
$i\succ_{p}j\quad\text{but}\quad f(i)\not\succ_{p^{\prime}}f(j),$
@@
Because both proposed base score extraction functions $\nu_{1}$ and $\nu_{2}$ are monotonic with respect to the preference order, a change in the ordering of any pair of arguments necessarily alters their relative distances $d_{i},d_{j}$ and hence their base scores $\tau(i)$ and $\tau(j)$. Consequently,
$\nu_{p}(X)\neq\nu_{p^{\prime}}(X),$
@@
$\nu_{1}(1,0,D,d)\;=\;0+(1-0)\,\frac{D-d}{D-1}\;=\;\frac{D-d}{D-1}.$
For the most preferred argument, we have $d=1$, hence
$\nu_{1}(1,0,D,1)\;=\;\frac{D-1}{D-1}\;=\;1.$
For the least preferred argument, we have $d=D$, hence
$\nu_{1}(1,0,D,D)\;=\;\frac{D-D}{D-1}\;=\;0.$
Because $d\in\{1,\dots,D\}$ and the function (Dd)/(D1)(D-d)/(D-1) decreases linearly from $1$ to 0 as $d$ goes from $1$ to $D$, all base scores lie in $[0,1]$ with the required endpoints. Therefore, $\nu_{1}$ with $\top=1$ and $\bot=0$ satisfies the Base Score Normalisation property. ∎
@@
$\nu_{1}(1-\bot,\bot,D,d)\;=\;\bot+(1-2\bot)\,\frac{D-d}{D-1}.$
For the most preferred argument ($d=1$),
$\nu_{1}(1-\bot,\bot,D,1)\;=\;\bot+(1-2\bot)\,\frac{D-1}{D-1}=\bot+(1-2\bot)=1-\bot=\top.$
For the least preferred argument ($d=D$),
$\nu_{1}(1-\bot,\bot,D,D)\;=\;\bot+(1-2\bot)\,\frac{D-D}{D-1}=\bot.$
@@
$\nu_{2}(0,0,D,d)\;=\;\frac{D-d}{D-1}.$
For the most preferred argument, we have $d=1$, hence
$\nu_{2}(0,0,D,1)\;=\;\frac{D-1}{D-1}\;=\;1.$
For the least preferred argument, we have $d=D$, hence
$\nu_{2}(0,0,D,D)\;=\;\frac{D-D}{D-1}\;=\;0.$
Since $d\in\{1,\dots,D\}$, the value (Dd)/(D1)(D-d)/(D-1) ranges linearly from $1$ to 0 as $d$ goes from $1$ to $D$, and all values lie in $[0,1]$. Therefore, $\nu_{2}$ with $\alpha=\beta=0$ satisfies the Base Score Normalisation property. ∎
@@
Given the definition of $\nu_{2}$, for the most preferred argument ($d=1$) we obtain the upper limit
$\top\;=\;\nu_{2}(\alpha,\beta,D,1)=\frac{D-1+\alpha}{D-1+\beta},$
and for the least preferred argument ($d=D$), the lower limit is
$\bot\;=\;\nu_{2}(\alpha,\beta,D,D)=\frac{\alpha}{D-1+\beta}.$
@@
i.e. $\alpha=\beta/2$. Hence, setting $\alpha=\beta/2$ ensures $\top=1$-$\bot$, so $\nu_{2}$ satisfies the Base Score Centralisation property.
@@
$\tau(x)-\tau(y)$ $=\nu_{1}(\top,\bot,D,d_{x})-\nu_{1}(\top,\bot,D,d_{y})$
$=(\top-\bot)\frac{(D-d_{x})-(D-d_{y})}{D-1}$
$=(\top-\bot)\frac{d_{y}-d_{x}}{D-1}.$
Hence if $d_{y}-d_{x}$ is the same (e.g. $\delta$ or $\Delta$) for both pairs, then $\tau(a)-\tau(b)=\tau(c)-\tau(d)$. For the indifference relation ($\simeq$), $d_{y}-d_{x}=0$ and the difference is 0.
@@
$\nu_{2}(\alpha,\beta,D,d)=\frac{D-d+\alpha}{D-1+\beta}.$
@@
$\tau(x)-\tau(y)=\frac{D-d_{x}+\alpha}{D-1+\beta}-\frac{D-d_{y}+\alpha}{D-1+\beta}=\frac{d_{y}-d_{x}}{D-1+\beta}.$
Again, this difference depends only on $d_{y}-d_{x}$; therefore, equal increments $\delta$ or $\Delta$ produce equal base score differences for the two pairs, and indifference yields zero difference.
@@
Let $\succeq_{p}$ be a preference ordering over $\mathcal{X}$ and let $S$ be a set of new arguments such that $\mathcal{X}^{\prime}=\mathcal{X}\cup S$ and for every $s\in S$ there exists $a\in\mathcal{X}$ with $a\simeq s$ under the extended ordering $\succeq_{p^{\prime}}$. By construction, adding $S$ introduces only indifference relations ($\simeq$) between new arguments and some existing arguments in $\mathcal{X}$. Observe two points about Algorithm 1 and the distance assignment (Definition 9):
@@
    The total distance span $D=1+n\cdot\delta+m\cdot\Delta$ depends only on the counts $n$ and $m$ of $\succ$ and $\succ\!\succ$ relations in the ordering. Since every $s\in S$ is indifferent to some $a\in\mathcal{X}$, the numbers $n$ and $m$ for relations between distinct preference tiers remain unchanged; hence $D$ is unchanged.
@@
    When traversing the ordering, Algorithm 1 increases the running distance $d$ only when the previous relation was $\succ$ (by $\delta$) or $\succ\!\succ$ (by $\Delta$), and leaves $d$ unchanged when the previous relation is $\simeq$. Therefore, inserting arguments that are indifferent to existing ones does not change the distance values assigned to the original arguments in $\mathcal{X}$ (they either share the same distance as before or receive the same $d$ as their indifferent partner).
Because both $\nu_{1}(\top,\bot,D,d)$ and $\nu_{2}(\alpha,\beta,D,d)$ depend only on $D$ and the per-argument distances $d$, and neither $D$ nor the distances of original arguments change under the described insertion of indifferent arguments, it follows that the base scores assigned to arguments in $X$ remain identical:
$\nu_{p}(\mathcal{X})\;=\;\nu_{p^{\prime}}(\mathcal{X}^{\prime}\setminus S).$

### v3

@@ 5. Evaluation
This section presents a theoretical validation of the concrete BSEFs based on how they satisfy the theoretical properties we defined. We also undertake an experimental evaluation, showing the pros and cons of each BSEF. Finally, we give an intuition of which gradual semantics to select depending on the decision-making context.
@@ 5.1. Theoretical Analysis
First, we show under which conditions our proposed BSEFs satisfy the axioms and properties presented. This is summarised in Table 1. The proofs can be found in Civit_arxiv26.
@@
Setting the constants $\Delta>0$ and $\delta>0$ satisfies Axiom 1 for both BSEFs.
Preference ordering $\succeq$ Design Choices QE Strengths EB Strengths DF Strengths
@@
Table 2. Examples of preference orderings for the running example with different design choices for the BSEF $\nu_{1}$ (Eq. 1). The value in bold between $\sigma_{slow}$ and $\sigma_{fast}$ is the system output. The (-) symbol indicates when a design choice is not needed, which is the case when all arguments are equally preferred, and their base scores are set to 0.5.
@@
Setting $\Delta>\delta>0$ satisfies Axiom 2 for both BSEFs.
Since the BSEFs are based on the distance between arguments given an ordering, if that ordering is changed, the base scores’ ordering will also change.
@@
Both BSEFs satisfy Axiom 3.
@@
To satisfy Property 3 (Base Score Relation Regularity), the algorithm variables to compute the distance ($\Delta,\delta$) must be constant.
@@
Both BSEFs satisfy Base Score Preference Stability (Property 4).
@@
Given a BSEF $\nu$, where $\nu$ satisfies Axiom 1, and a semantics $\sigma$, where $\sigma$ satisfies monotonicity and balance, for any arguments $a$, $b$ where $a\succ b$ and $a$ and $b$ have equivalent supporters and attackers, then $\sigma(a)>\sigma(b)$.
@@ 5.2. Experimentation
To explore the robustness of our approach, we generated 30,000 random samples of preference orderings and design choices, recording the option selected by three gradual semantics. These semantics are: QE Model, Euler-Based (EB) Amgoud_IJAR18, and DF-QuAD (DF) Rago_kr16, which are among the most used in the literature. The decision arguments’ base scores are set to 0.5. We computed pairwise agreement and Cohen’s Kappa coefficient ($\kappa$) to assess consistency between semantics. Under the centralisation property, the pairwise agreement values were QEEB=0.98QE-EB=0.98, QEDF=0.85QE-DF=0.85, and EBDF=0.84EB-DF=0.84. The corresponding Cohen’s Kappa scores were QEEB=0.96QE-EB=0.96, QEDF=0.65QE-DF=0.65, and EBDF=0.62EB-DF=0.62. Without centralisation, the pairwise agreement values were QEEB=0.96QE-EB=0.96, QEDF=0.81QE-DF=0.81, and EBDF=0.81EB-DF=0.81, with Cohen’s Kappa scores of QEEB=0.92QE-EB=0.92, QEDF=0.52QE-DF=0.52, and EBDF=0.52EB-DF=0.52. These results show a strong alignment between QE and EB, while DF-QuAD diverges more, particularly when base score ranges are extreme (e.g., $\tau(e)\approx 1$, $\tau(e)\approx 0$). This difference shows the structural properties of the methods discussed in Sec. 5.3.
To further analyse the influence of user preferences and design parameters, Table 2 reports examples of preference orderings combined with different design choices, evaluated under the three gradual semantics. The Base Score Relation Regularity property is assumed to reduce experimental complexity and noise.
@@ 5.3. Discussion on the Gradual Semantics
Figure 2. Influence of an argument’s base score and its supporters or attackers on the final argument strength under different gradual semantics. Each subplot corresponds to a distinct semantics: DF-QuAD (left), Quadratic Energy (centre), and the Euler-based Semantics (right). The x-axis represents the base score of the influenced argument, and the y-axis its resulting final strength. Blue lines denote when the influencing argument is a supporter (+), while red lines denote when it is an attacker (-). Each line has assigned its aggregation ($\alpha$), with colour intensity indicating its strength.
@@
Figure 2 shows the influence of a single attacker or supporter (the influencer), and an argument’s base score (the influenced) for its final strength. The DF-QuAD semantics use product aggregation with a linear influence function, which causes the influencer’s base score to dominate the final outcome. When an attacking (or supporting) argument has a base score close to 1, it almost completely drives the influenced argument’s strength toward 0 (or 1), regardless of the influenced base score. This strong reactivity can be desirable in safety-critical scenarios, where highly preferred arguments should decisively determine the outcome, but it may reduce robustness in settings with uncertain or noisy preferences. In contrast, the QE semantics combines sum aggregation with a 2-Max influence function. Here, both the influencer and the influenced arguments contribute comparably to the final strength, leading to smoother transitions and more gradual adaptation to preference changes (e.g., given an influencer with a base score of 1, the influenced’s final strength will be $\sigma=0.5\tau+0.5$ if it receives a support, and $\sigma=0.5\tau$ if it receives an attack). QE, therefore, offers a good compromise between responsiveness and stability, which makes it particularly suitable for explainable decision-support systems that require consistent reasoning traces. Finally, the EB semantics is the most conservative, using sum aggregation with an Euler-based influence function. When an argument’s base score is close to the extremes (0 or 1), its final strength remains nearly fixed at those values, regardless of new supports or attacks. This behaviour preserves high-confidence arguments, making EB appropriate when users insist on arguments being non-negotiable or when decisions must respect rigid value priorities.
@@ 6. Related Work
Gradual argumentation relies on the arguments’ base scores, and its literature has unveiled different ways of determining them depending on the context. In Baroni_AC15, the QuAD framework was introduced to support debates on design alternatives. Since these scenarios do not involve a large number of voters, base scores are assessed by experts or derived from predefined criteria. As an extension, in Rago_icppmas17, the authors presented QuAD-V, which allows users to vote for or against arguments, using these votes to compute base scores. A similar procedure is carried out in Tarle_aamas22. In another work Cocarascu_aamas19 in which a QBAF is generated from movie reviews, the base scores are obtained from an aggregation of critics’ votes. Another use of QBAFs is to apply their methods for explainable AI, e.g. by transforming (deep) neural networks into QBAFs Albini_arxiv20; DBLP:conf/ecai/AyoobiPT23. For instance, in Albini_arxiv20, the base scores of the arguments are obtained from the attacking, supporting neurons, and their activations in the network, similarly to Potyka_aaai21. In contrast, the work from Battaglia_icsum24 sets base scores according to users’ preferences to personalise decision support, similarly to our work. There, base scores are initially set arbitrarily and then discounted for less preferred arguments. This work similarly relies on preferences, but uses them to determine the initial base scores directly.
Within structured argumentation, the work in Tamani_icfs14 focuses on ASPIC argumentation frameworks with fuzzy set theory, using expert knowledge to assess argument importance, similarly to a base score.
In summary, two main limitations are found in existing work. First, most approaches do not account for the preferences of a single user when setting base scores, limiting personalisation in argumentation-based decision-making systems. Second, when preferences are considered, base scores are initialised arbitrarily and adjusted afterwards. In this work, we address those research gaps.
@@ 7. Conclusions and Future Work
This work introduces BSEFs, a principled method to derive argument base scores from user preference orderings in gradual argumentation. We defined a set of axioms and desirable properties, proposed two concrete BSEFs with tunable design choices, and incorporated non-linear preference intensity to better capture human reasoning. Theoretical analysis confirmed that the functions satisfy coherence and regularity properties, while experiments in a robotics scenario showed that preference ordering dominated the decision outcome and design choices primarily modulate sensitivity.
Future directions include the fine-tuning or definition of the different design choices given the context and user feedback, allowing partial orderings in preferences in case some arguments cannot be compared Visser_wtafa11, and computing the aggregation when using this approach when considering more than one user’s preferences in the decision-making Rago_AI25. A philosophical question which should be commented on is whether it is realistic to consider the scenario in which the most preferred argument supports the least preferred argument. Our focus was on bipolar argumentation, but gradual structured argumentation frameworks are receiving attention lately Rago_kr25; Rapberger_kr25. A future extension of this work is to adapt our method to structured argumentation. Furthermore, we plan to validate our approach by performing a user study. Finally, the extension of this work into value-based argumentation could provide an approach for more value-aligned decision-making Atkinson_AL21; Bodanza_AC23. This could be potentially achieved by ordering the values promoted by the arguments and assigning a base score to each value.
{acks}
This work was supported by the project CHLOE-MAP PID2023-152259OB-I00 funded by MCIU/ AEI /10.13039/501100011033 and by ERDF, UE; the project ROBOCAT SDC006/25/000016 funded by the Generalitat de Catalunya, NextGenerationEU. A. Civit has been supported by AGAUR-FI ajuts (2023 FI-3 00065) Joan Oró of the Generalitat of Catalonia and the European Social Plus Fund. F. Toni has been supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101020934).
@@ Appendix A Proofs
Let $a,b\in\mathcal{X}$ be two adjacent arguments in the ordering, such that $a\succ b$ in the preference ordering. According to Definition 9 and Algorithm 1, the distance assigned to $a$ and $b$ satisfies $d(b)=d(a)+\delta$. Where $\delta>0$ for a regular preference ($\succ$) and $\Delta$¿0 for a much greater preference ($\succ\!\succ$). Both proposed base score extraction functions $\nu_{1}$ and $\nu_{2}$ are strictly decreasing with respect to $d(x)$. Therefore, if $d(a)<d(b)$, it follows that $\nu(\succeq)(a)>\nu(\succeq)(b)$. This implies that the base score of the preferred argument $a$ is greater than that of $b$, i.e., $\tau(a)>\tau(b)$. Hence, setting $\delta>0$ (and similarly $\Delta>0$) ensures that both base score extraction functions satisfy Axiom 1. ∎
@@
$d(b)=d(a)+\Delta,\qquad d(c)=d(b)+\delta,\qquad d(d)=d(c),$
@@
which satisfies Axiom 2. Therefore, setting $\Delta>\delta>0$ guarantees that both $\nu_{1}$ and $\nu_{2}$ respect the intended magnitude differences between much greater and greater preferences. ∎
@@
Let $\nu(\succeq)$ be a base score extraction function applied to a set of arguments $X$ with a preference ordering $\succeq$. By definition, $\nu(\succeq)$ assigns base scores according to the relative ordering and distances among the arguments.
Suppose now that the preference ordering changes to a non-isomorphic ordering $\succeq^{\prime}$. By Definition 1, two preference orderings are isomorphic if and only if there exists a bijection $f:X\to X^{\prime}$ such that, for all $i,j\in X$,
$i\succeq j\;\Leftrightarrow\;f(i)\succeq^{\prime}f(j).$
If $\succeq$ and $\succeq^{\prime}$ are not isomorphic, then at least one pair of arguments $(i,j)$ exists such that their preference relations differ, i.e.,
$i\succ j\quad\text{but}\quad f(i)\not\succ^{\prime}f(j),$
@@
Because both proposed base score extraction functions $\nu_{1}$ and $\nu_{2}$ are monotonic with respect to the preference order, a change in the ordering of any pair of arguments necessarily alters their relative distances $d(i)$, $d(j)$ and hence their base scores $\tau(i)$ and $\tau(j)$. Consequently,
$\nu(\succeq)\neq\nu(\succeq^{\prime}),$
@@
$\nu_{1}(\succeq)(x)\;=\;0+(1-0)\,\frac{D-d(x)}{D-1}\;=\;\frac{D-d(x)}{D-1}.$
For the most preferred arguments $M\in\mathcal{X}$, we have $\forall m\in M,\ d(m)=1$, hence
$\nu_{1}(\succeq)(m)\;=\;\frac{D-1}{D-1}\;=\;1.$
For the least preferred arguments $L\in\mathcal{X}$, we have $\forall l\in L,\ d(l)=D$, hence
$\nu_{1}(\succeq)(l)\;=\;\frac{D-D}{D-1}\;=\;0.$
Because $d(x)\in\{1,\dots,D\}$ and the function (Dd(x))/(D1)(D-d(x))/(D-1) decreases linearly from $1$ to 0 as $d(x)$ goes from $1$ to $D$, all base scores lie in $[0,1]$ with the required endpoints. Therefore, $\nu_{1}$ with $\top=1$ and $\bot=0$ satisfies the Base Score Normalisation property. ∎
@@
$\nu_{1}(\succeq)(x)\;=\;\bot+(1-2\bot)\,\frac{D-d(x)}{D-1}.$
For the most preferred arguments ($m\in\mathcal{X}$ such that $d(m)=1$),
$\nu_{1}(\succeq)(m)\;=\;\bot+(1-2\bot)\,\frac{D-1}{D-1}=\bot+(1-2\bot)=1-\bot=\top.$
For the least preferred arguments ($l\in\mathcal{X}$ such that $d=D$),
$\nu_{1}(\succeq)(l)\;=\;\bot+(1-2\bot)\,\frac{D-D}{D-1}=\bot.$
@@
$\nu_{2}(\succeq)(x)\;=\;\frac{D-d(x)}{D-1}.$
For the most preferred arguments $m\in\mathcal{X}$, we have $d(m)=1$, hence
$\nu_{2}(\succeq)(m)\;=\;\frac{D-1}{D-1}\;=\;1.$
For the least preferred arguments $l\in\mathcal{X}$, we have $d(l)=D$, hence
$\nu_{2}(\succeq)(l)\;=\;\frac{D-D}{D-1}\;=\;0.$
Since $d(x)\in\{1,\dots,D\}$, the value (Dd(x))/(D1)(D-d(x))/(D-1) ranges linearly from $1$ to 0 as $d(x)$ goes from $1$ to $D$, and all values lie in $[0,1]$. Therefore, $\nu_{2}$ with $\alpha=\beta=0$ satisfies the Base Score Normalisation property. ∎
@@
Given the definition of $\nu_{2}$, for the most preferred arguments $m\in\mathcal{X}$ ($d(m)=1$) we obtain the upper limit
$\top\;=\;\nu_{2}(\succeq)(m)=\frac{D-1+\alpha}{D-1+\beta},$
and for the least preferred arguments $l\in\mathcal{X}$ ($d(l)=D$), the lower limit is
$\bot\;=\;\nu_{2}(\succeq)(l)=\frac{\alpha}{D-1+\beta}.$
@@
i.e. $\alpha=\beta/2$. Hence, setting $\alpha=\beta/2$ ensures $\top=1$-$\bot$, so $\nu_{2}$ satisfies the Base Score Centralisation property. ∎
@@
$\tau(x)-\tau(y)$ $=\nu_{1}(\succeq)(x)-\nu_{1}(\succeq)(y)$
$=(\top-\bot)\frac{(D-d(x))-(D-d(y))}{D-1}$
$=(\top-\bot)\frac{d(y)-d(x)}{D-1}.$
Hence if d(y)d(x)d(y)-d(x) is the same (e.g. $\delta$ or $\Delta$) for both pairs, then $\tau(a)-\tau(b)=\tau(c)-\tau(d)$. For the indifference relation ($\simeq$), d(y)d(x)=0d(y)-d(x)=0 and the difference is 0.
@@
$\nu_{2}(\succeq)(x)=\frac{D-d(x)+\alpha}{D-1+\beta}.$
@@
$\tau(x)-\tau(y)=\frac{D-d(x)+\alpha}{D-1+\beta}-\frac{D-d(y)+\alpha}{D-1+\beta}=\frac{d(y)-d(x)}{D-1+\beta}.$
Again, this difference depends only on d(y)d(x)d(y)-d(x); therefore, equal increments $\delta$ or $\Delta$ produce equal base score differences for the two pairs, and indifference yields zero difference.
@@
Let $\succeq$ be a preference ordering over $\mathcal{X}$ and let $S$ be a set of new arguments such that $\mathcal{X}^{\prime}=\mathcal{X}\cup S$ and for every $s\in S$ there exists $a\in\mathcal{X}$ with $a\simeq s$ under the extended ordering $\succeq^{\prime}$. By construction, adding $S$ introduces only indifference relations ($\simeq$) between new arguments and some existing arguments in $\mathcal{X}$. Observe two points about Algorithm 1 and the distance assignment (Definition 9):
@@
    The maximum distance span $D=1+n\cdot\delta+m\cdot\Delta$ depends only on the counts $n$ and $m$ of $\succ$ and $\succ\!\succ$ relations in the ordering. Since every $s\in S$ is indifferent to some $a\in\mathcal{X}$, the numbers $n$ and $m$ for relations between distinct preference tiers remain unchanged; hence $D$ is unchanged.
@@
    When traversing the ordering, Algorithm 1 increases the running distance $d(x)$ only when the previous relation was $\succ$ (by $\delta$) or $\succ\!\succ$ (by $\Delta$), and leaves $d(x)$ unchanged when the previous relation is $\simeq$. Therefore, inserting arguments that are indifferent to existing ones does not change the distance values assigned to the original arguments in $\mathcal{X}$ (they either share the same distance as before or receive the same $d(x)$ as their indifferent partner).
Because both $\nu_{1}(\succeq)(x)$ and $\nu_{2}(\succeq)(x)$ depend only on $D$ and the per-argument distances $d(x)$, and neither $D$ nor the distances of original arguments change under the described insertion of indifferent arguments, it follows that the base scores assigned to arguments in $X$ remain identical:
$\nu(\succeq)\;=\;\nu^{\prime}(\succeq^{\prime}).$